### v1

# Environment-Specific People

Abstract
Despite significant progress in generative image synthesis and full-body generation in particular, state-of-the-art methods are either context-independent, overly reliant to text prompts, or bound to the curated training datasets, such as fashion images with monotonous backgrounds. Here, our goal is to generate people in clothing that is semantically appropriate for a given scene. To this end, we present ESP, a novel method for context-aware full-body generation, that enables photo-realistic inpainting of people into existing “in-the-wild” photographs. ESP is conditioned on a 2D pose and contextual cues that are extracted from the environment photograph and integrated into the generation process. Our models are trained on a dataset containing a set of in-the-wild photographs of people covering a wide range of different environments. The method is analyzed quantitatively and qualitatively, and we show that ESP outperforms state-of-the-art on the task of contextual full-body generation.
Figure 1: ESP generates people in semantically appropriate, environment-inspired clothing: given a 2D pose (row 1) and an environment (row 2); people are inpainted into the photo, based on a context-aware generative model that predicts contextual human parsing maps.
1 Introduction
In recent years, great progress has been made on all aspects of image generation and editing, e.g., GAN-based methods [53, 56, 49, 28, 54, 55] allow for photo-realistic image synthesis on well-crafted datasets, and diffusion-based models [41, 52, 32] make it possible for the original image content to be seamlessly removed or replaced via inpainting. Due to the rapid developments in generative AI, distinguishing between AI-generated content and real content by analyzing the quality of the generated images alone is getting more difficult each day. In this paper, we focus on full-body generation and inpainting of environment-specific people wearing clothing that is appropriate for a given environment. For example, if the context of the depicted scene is a football field, our method would generate a person dressed in sportswear such as shorts, a T-shirt or a polo shirt, football boots, etc. However, if the context switches to snowy mountains, the inserted person would be wearing a winter jacket, ice skates, or ski attire.
One line of work in full-body generation is centered around using StyleGAN [11, 12] to generate full-body samples. Such methods are usually trained on high-quality data, such as DeepFashion [31], that is curated and hand annotated. To reduce its high dimensionality, data is first normalized and filtered. This is a crucial step that is contributing to the success of StyleGAN-based methods. While the images that are generated in this way are of high-quality, the main issue that we tackle in this paper is the semantic asyncrony between the clothing and the corresponding environment that exists in these datasets due to the unnatural data acquisition settings. I.e., in fashion photography, images are usually captured in a studio setting, where the subject is fully lit and the background is very simple, e.g., a white screen. Therefore there is minimal environmental context available. Also, in case of an outdoor photo-shoot, the link between the environment and the clothing may be atypical and it does not represent what is normally worn on the location well.
The other, more recent line of work, are diffusion-based methods or, typically, large pre-trained latent diffusion models such as Stable Diffusion [41] and its conditional variations (e.g. ControlNet [52]). The main advantage of such methods is the existence of large pre-trained image priors that are usually fine-tuned on human data including full-bodies in the wild. However, in order to generate sufficiently varied samples, these txt2img models require accurate textual descriptions of the desired output. While providing such descriptions may lead to satisfactory results, in this work we assume that there are no textual descriptions of the appropriate clothing available. And instead, we focus on visual information alone, investigating whether the image context can be used as a sole driving source for the generation of environment-specific people.
To this end, we take advantage of both StyleGAN-based and diffusion-based methods, combining their strengths to devise a purely image-based system for our task. As part of our pipeline, contextual cues are first extracted from a given environment photo with a VAE and then fed into a StyleGAN-based human parsing map (HPM) generator that predicts context-aware clothing semantics. HPM training data is filtered and normalized to include only simple poses, making it suitable for training a StyleGAN model. Next, a HPM translation module is fine-tuned using the same dataset and in turn utilized to provide additional guidance to the inpainting process of a large pre-trained Stable Diffusion model, for seamless inpainting of photo-realistic and environment-specific people into the original photographs.
Our method is evaluated quantitatively and qualitatively, with a focus on contextual appropriateness. The results show that our method generates environment-specific people that match the context of the input photograph well, considering the clothing semantics (see Figure 1). Furthermore, our approach outperforms state of the art on the task of contextual full-body generation.
Our contributions are as follows:
  • a StyleGAN-based method for stochastic generation of context-aware HPMs that are generated using our newly introduced contextual style vectors;
  • a dynamic pix2pix-based extension of the computational graph that is fine-tuned in parallel with the StyleGAN generator which enables pose-guided imgage-to-image translation;
  • a system that leverages a pre-trained text-to-image latent diffusion model (Stable Diffusion), allowing for super-resolution imaging and photorealistic inpainting of environment-specific people into existing photographs.
Figure 2: System Overview: (I) an input context image is encoded into the latent space of a VAE, giving a contextual style vector; (II) the latter is then fed, alongside a random vector, into a pose-conditioned StyleGAN-II HPM generator; (III) an HPM translation module is used to achieve fine-grained control over the generated clothing during the inpainting process (using the generic Stable Diffusion LDM).
2 Related work
From their advent, Generative Adversarial Networks (GAN) [13] have proven to be a very effective tool for synthesizing realistic images. StyleGAN-based methods [20, 21, 23, 4] have been used to generate photo-realistic imagery of a variety of objects and scenes [50, 51] , including humans [26, 45]. More recently, large latent diffusion models have been in the spotlight [37] due to their unparalled image generation capabilities, conditioned on text (txt2img)  [41, 38, 39, 1, 42] and additional input controls  [52].
Full-body generation:
StyleGAN-based state-of-the-art for unconditional full-body generation include InsetGAN [11] and StyleGAN-Human [12]. Both methods use well prepared datasets that are either proprietary or contain fashion imagery. InsetGAN composes multiple GANs to handle the complex task of generating high resolution full body human images. One GAN is trained to generate the global structure, i.e., the body, while the remaining GANs are each trained separately for different local regions, e.g., face or shoes. The latent spaces of different GANs are jointly optimized for a seamless combination to synthesize the whole person. StyleGAN-Human involves training StyleGAN2 [22] on a large, self-curated dataset, consisting of around 230K full body human images. The background in this data is often white or plain, or the link between the background and the clothing is atypical. We use a small dataset of in-the-wild imagery [57] for training, where people are clothed synchronously with the environment, e.g. football gear is worn on a football field and warm winter attire is worn in the snow. Pose and/or appearance information are used to control the process of conditional full-body generation [3, 33, 34, 47, 44]. In a two step process, Lewis et al. [26] first train a dense-pose conditioned StyleGAN2 model with a clothing segmentation branch, after which a new image is synthesized using layered latent space interpolation to transfer a garment to a target body. Grigorev et al. [14] combine 3D parametric mesh-based models of human bodies [36] with neural textures generated by a modified StyleGAN2 model. Ren et al. [40] propose a deep architecture, conditioned on human pose, to combine flow-based operations with an attention mechanism. This architecture has also been extended for use in an unsupervised setting [43]. Sarkar et al. [45] warp encoded part-based latent appearance vectors in a normalized pose-independent space to generated synthetic people in specific poses. Our twist considering this line of work is to tackle the problem from a novel perspective, by incorporating image context as a conditioning through the mapping network. In this way, we generate environemnt-specific people that may be inpainted into the existing photographs, while respecting the scene context in terms of the clothing semantics.
Inpainting:
The task of generating new content in specified regions of an existing image is referred to as inpainting. This is a well-studied problem that is particularly suitable for our task [49, 28, 55]. Motivated by the StyleGAN2 modulated convolutions, Zhao et al. [54] improve inpainting reconstruction by using a comodulation layer. Approaches that build on a variational auto-encoder based architecture have also been proposed [56, 53]. Lugmayr et al. [32] improved on the previous methods by harnessing the expressiveness of a pre-trained denoising diffusion probabilistic model (DDPM) [17], which had previously shown impressive results in guided image synthesis in pixel space [9, 35]. Using a pre-trained DDPM as a generative prior, they alter the reverse diffusion iterations by sampling the unmasked regions using the given image information. However, this approach slows down the image generation process of a DDPM. More recently, similar ideas in the realm of latent diffusion models have been explored using denoising diffusion implicit models (DDIM) [48] where the denoising steps are skipped for faster inference. Most recent image generators that have the inpainting capacity use this principle and we take advantage of these models  [41].
Humans in scenes:
Brooks et al. [5] generate both the scene and a pose-compatible human that matches the scene context at 128x128, where the context embeddings are calculated from a 2D pose vector and a noise vector. In their work, the focus is on generating both compatible scenes and humans inside, conditioned on 2D pose embeddings. In contrast, our focus is on inpainting environment-specific people into existing scenes and we calculate context embeddings from the scene context directly. Kulal et al. [25] filter a large-scale video dataset (2.4 million clips) at 256x256 and use it to train a multi-faceted LDM, generating pose-compatible people conditioned on a photo of a reference person and a masked scene. The primary focus of this work is on synthesizing plausible poses given a scene context. For this purpose, two random frames that contain the same human from each video are used during training. At test time, the reference person conditioning may be removed, allowing for the synthesis of scene-compatible humans, given scene information. In contrast, we condition the generation process on a full-body 2D pose rendering (similar to  [8]) and investigate the connection between the scene and its impact on clothing alone, given a fixed 2D pose. Additionally, we only use a small dataset containing approx. 23K images at 200x200, and our results are generated at 512x512. Image quality of our method is on par with state-of-the-art image generators including ControlNet  [52] and Stable Diffusion [41] that many other recent high-quality full-body generation methods build on top of  [30].
3 Method
Our system is shown in Figure 2. First, context embeddings are extracted from image $\mathcal{I}$ that contains the environmental context $\mathcal{C}$. Then, using the extracted embeddings, small human parsing maps $\mathcal{M}$ are generated conditioned on a 2D skeleton rendering $\mathcal{S}$ (32x64). Next, generated HPMs are then used, with the help of our HPM translation module, as an additional guiding signal for the seamless inpainting of environment-specific people into the existing photographs. In this final step, we take advantage of a large pre-trained LDM (Stable Diffusion for inpainting), allowing for the generation of images at very high resolution (512x512).
3.1 Context embeddings
We must first find a representation of $\mathcal{I}$ that can be used as the context conditioning to drive the generation of the clothing semantics in the following steps. During training, we have access to the context image containing the human $I_{c}^{h}$ and a corresponding mask of the human $M_{h}$. We dilate the mask with morphological operations and invert it to get the soft background mask $M_{b}$. To remove the foreground human, we apply a pixel-wise mask on the input image, producing the masked image $\mathcal{I}=I_{c}^{h}\odot M_{b}$ Variational auto-encoders (VAE) [24] are good at representing the overall global information present in an image [15], so we choose the compact and normalized latent space of a VAE to reflect the global structure, illumination and color tone of the context image $\mathcal{I}$. We train the VAE with $i_{c}$ where $i_{c}$ is a down-sampled $64\times 64$ version of $\mathcal{I}$ with the loss:
$\begin{split}L_{V}=\lambda_{1}\ ||i_{c}-\tilde{i}_{c}||_{1}+\lambda_{2}D_{KL}({\scriptstyle\mathcal{N}(\mu_{c},\sigma^{2}_{c})}||{\scriptstyle\mathcal{N}(0,1)}),\end{split}$ (1)
where $\tilde{i}_{c}$ is the reconstructed image from the VAE; $\lambda$’s are the corresponding weights for the different losses; $D_{KL}$ is the KL divergence loss [24]; $\mathcal{N}(x,y)$ is the normal distribution with mean $x$ and variance $y$; and $\mu_{c}$, $\sigma^{2}_{c}$ are the mean and variance of the context image. Once trained, the latent vector $\mathcal{C}$ sampled from the VAE for corresponding context image $\mathcal{I}$ is used for conditioning in the following steps. Note that the reconstruction loss term is applied only to the regions where humans are present and as such we do not require $\mathcal{I}$ to be computed from $I_{c}^{h}$ at inference time. Additionally, $\tilde{i}_{c}$ is subsequently passed through the VAE to remove the small artifacts that may appear in the regions that people typically occupy in our training data.
3.2 Contextualized generation of HPMs
We train a StyleGAN2 generator, conditioned on $\mathcal{C}$, to synthesize a human parsing map (HPM) and a low-resolution reconstruction of the corresponding context image, for a given 2D pose rendering $\mathcal{S}$. A 512-dimensional random vector $z_{s}$, sampled from $\mathcal{N}(0,1)$, is concatenated with $\mathcal{C}$ (also 512-d) and then passed as input to the StyleGAN2 [22] generator. After $z_{s}$ and $\mathcal{C}$ are passed through the mapping network, they are jointly referred to as the contextual style vector in this paper. Concatenating $z_{s}$ with $\mathcal{C}$ is important because there may be more than one outfit that is appropriate for a given environment. $z_{s}$ captures the one to many mapping between the context and the semantics.
The HPM generator is additionally conditioned with $\mathcal{S}$ in the image space to enforce the generation of full-body humans in the given pose. This helps us isolate the pose later on in the experiments and focus on exploring the impact that the environment by itself has on the clothing semantics. Motivated by Isola et al. [19], to inject the pose conditioning into the generator, we concatenate the input of different style blocks of the generator with the output of the corresponding layers of a pose encoder. The input to the encoder is an RGB image $\mathcal{S}$ and its layers are designed so that the number of the feature maps and their dimensions match the StyleGAN2 generator (32x64). The generator and the pose encoder are trained end-to-end.
The HPM generator jointly outputs a HPM $\tilde{\mathcal{M}}$ (3 channels) and a blurry background image (3 channels). Note that the blurry background reconstruction from the VAE that is used as the GT image for the generator is rescaled from 64x64 to 32x64 to match the HPM image dimensions. Similar to [11], we observe that GANs do not handle large areas of uniform color well and therefore we use the tall HPM images where the number of the background pixels has been reduced, rather than square images with many background pixels. We are able to do so because our dataset is filtered and it only contains the poses where the width of the person is equal to at most one half of their height.
Finally, to help the generator understand the link between the clothing semantics and the context, we use as a joint (HPM, background) discriminator on top of the individual HPM and background discriminators which are used to increase the quality and realism of generated HPMs and backgrounds, respectively. In summary, the generator loss for the semantic maps is defined as:
$\begin{split}L_{S}^{G}=L_{adv}^{D_{\mathcal{M}}}+L_{adv}^{D_{B}}+L_{adv}^{D_{MB}}\\ +\lambda_{3}L_{1}+\lambda_{4}L_{R}+\lambda_{5}L_{perc}\end{split}$ (2)
where $L_{adv}^{D_{\mathcal{M}}}$ and $L_{adv}^{D_{B}}$ represent the adversarial losses that come from the HPM and the blurry background image discriminators respectively; $L_{adv}^{D_{MB}}$ is the loss from an additional discriminator that evaluates the composite of the blurry background and the HPM; $L_{1}$ is the $l_{1}$ loss applied between the generated blurry background and the corresponding generated background $\tilde{i}_{c}$ from the VAE; $L_{perc}$ is a standard perceptual loss commonly used in the literature for generating more detailed images; $L_{R}$ represents the path regularisation [22] that is applied to the generator; $\lambda_{3}$, $\lambda_{4}$, and $\lambda_{5}$ are the weighting factors for the specific losses.
3.3 HPM translation module
In the remainder of the pipeline, we use diffusion models [17]. In particular, we leverage the large pre-trained latent diffusion model that is called Stable Diffusion [41] as our image prior. Diffusion models are well-known for their capacity to generate highly detailed images. They do so gradually through the denoising diffusion process in $T$ timesteps. During training, Gaussian noise $\mathcal{N}(\mathbf{0},\mathbf{I})$ gets added to the original sample $\mathbf{x}_{0}$, step by step, and a model similar to UNet [18] is trained to predict the noise $\epsilon$ residuals, at timestep $t$. For latent diffusion models where the denoising process is performed in the latent space rather than in pixel space, $\mathbf{z}_{t}$ marks the corresponding latent representation of a noisy sample $\mathbf{x}_{t}$ at timestep $t$. The LDM objective is then formulated as follows:
$L:=\mathbb{E}_{\mathcal{E}(x),\epsilon\sim\mathcal{N}(0,1),t}\Big{[}\|\epsilon-\epsilon_{\theta}(z_{t},t)\|_{2}^{2}\Big{]}\,.$ (3)
Given the large success of LDMs, new mechanisms have appeared on how to control their image generation process further. In this paper we use the approach called ControlNet from [52]. The main idea behind ControlNet is to create a trainable copy of a pre-trained LDM, attach it to the LDM and in turn train it without modifying the original model (zero-convolutions). In this way, generated images retain the quality from the original LDM which remains unchanged. The copy is additionally conditioned on images $\mathbf{c}_{\text{f}}$ that are encoded into its latent space by a small learnt neural network and the feature maps of the encoder of the LDM copy are added to the feature maps of the original LDM so as to transfer the original conditioning signal. Including the text conditioning $\mathbf{c}_{t}$, objective fuction for a conditional LDM is now equal to:
$\vspace{-3pt}\begin{split}L^{C}=\mathbb{E}_{\mathbf{z}_{0},\mathbf{t},\mathbf{c}_{t},\mathbf{c}_{\text{f}},\epsilon\sim\mathcal{N}(0,1)}\Big{[}\|\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},\mathbf{t},\mathbf{c}_{t},\mathbf{c}_{\text{f}}))\|_{2}^{2}\Big{]}\,.\end{split}$ (4)
We attach a ControlNet module to the Stable Diffusion model and train it on approx. 23K images where the input conditioning are RGB HPM images while the output are the corresponding images of real people in the wild. This model is later used as a translator for our custom low resolution RGB HPM images that are padded and then upscaled from 32x64 to 512x512 during training. The output is upscaled from 200x200 to 512x512 during training.
3.4 Inpainting and super-resolution
Human body is highly complex due to its high flexibility that stems from the many degrees of freedom that it has. Also, one body may be clothed in many different types of outfits. For this reason, there exists a large semantic gap between the body and its binary segmentation mask. This gap makes it especially difficult to inpaint human bodies into photographs, when only the ambiguous binary mask is available and the body is to be inpainted into the specified region. To this end, at inference time, we condition the inpainting LDM on a semantic representation of the body and clothing (HPM) via our pre-trained HPM translation module (ControlNet). In this way, we are able to transfer the background from the original image into the resulting image using the inpainting model while predicting the HPM-compatible humans at the same time. Additionally, by using a weighting scheme where the impact of the HPM translator is reduced, it is possible to generate images of very high quality where the quality is similar to that of the Stable Diffusion-generated imagery. This is done by reducing the weight $\beta$ (controlNet strength) that is used for the multiplicaton of the ControlNet’s feature maps prior to their addition with the feature maps of the inpainting model. However, $\beta$ should not be reduced drastically as the benefits of the guidance signal will become undone. We find that setting $\beta$ to $0.5$ works well.
Figure 3: Fixing the context (rows F) generates more uniform clothing than varying the context embeddings (rows R). This indicates there is a link between the context and the environment that has been learned by our context-aware HPM generator.
Figure 4: Background Reconstruction: VAE reconstructions are shown in rows 1 (GT) and 2 (VAE); fixing the context vector in the StyleGAN generator gives matching results (row F) with GT/VAE, while varying the context leads to random predictions (row R).
Figure 5: SOTA Comparison: the environment where the humans are to be inpainted is shown in the first row; results with StyleGAN-Human [12] are in row 2; ControlNet [52] results are in row 3. Our method generates more appropriate, environment-specific, people.
Output $\psi=0.5$ $\psi=0.6$ $\psi=0.7$ $\psi=0.8$ $\psi=0.9$ $\psi=1$
HPM 15.73 14.59 13.53 12.64 11.96 11.49
Context 112.45 86.33 67.55 54.77 46.29 41.08
Table 1: Truncation Trick: the table shows FIDnormal-$\downarrow$ scores for varying $\psi$ across 10K randomly selected contexts. The optimal value of $\psi$ is equal to $1.0$ for both the HPM and the Context generator.
Method $\beta=0$ $\beta=0.25$ $\beta=0.5$ $\beta=0.75$ $\beta=1$
Baseline - - 78.7 86.7 90.0
Ours 42.3 22.6 21.0 19.8 20.2
Table 2: ControlNet Strength: the table shows FIDnormal-$\downarrow$ scores for varying $\beta$ across 10K images. Our method outperforms SOTA.
4 Experiments
4.1 Data
Our method utilizes a training dataset comprising context images, images with humans, skeletal pose information, and machine-annotated human body parsing or semantic segmentation maps. For training with the AHP dataset [57], we extract context images by using segmentation masks to remove people from the images. Skeleton images with pose data are generated using OpenPose [7, 6] on these human-included context images. During inference, only a context image and pose information are required to generate a full-body person image with context-aware appearance.
C Evaluation PSNR $\uparrow$ MSE $\downarrow$ LPIPS $\downarrow$ FID $\downarrow$ KID $\downarrow$
HPM 20.44 0.0102 0.051 11.5 0.0098
F Blurry Context 23.33 0.0060 0.123 41.1 0.0412
Real Context 21.94 0.0078 0.154 - -
HPM 20.37 0.0104 0.053 12.2 0.0104
R Blurry Context 11.81 0.0769 0.361 41.4 0.0389
Real Context 11.71 0.0781 0.358 - -
F VAE Recon. 27.05 0.0022 0.051 7.0 0.0029
Table 3: Quantitative Analysis: the results are calculated over 10K contexts. C denotes whether the context vector is fixed (F) or random (R). Overall, fixing the context vector leads to the predictions that are closer to the ground truth, especially in background reconstruction. Therefore, the analysis confirms that our method is context-aware (PSNR, MSE, LPIPS) and that it generates meaningful parsing masks and background contexts (FID, KID). The performance of the standalone VAE module used to generate the blurry contexts (rows 2 and 5) is shown in the last row.
The AHP dataset is compiled from large-scale instance segmentation and detection datasets, including COCO [29], VOC [10], SBD [16], and Objects365 [46]. It underwent filtering and post-processing to yield 56,599 images with parsing maps and binary human masks, which are integral to our method. However, the raw AHP dataset’s varied resolution and quality, along with inconsistent human pixel percentages, lead to suboptimal results. Additionally, the provided parsing maps sometimes have disconnections and holes.
To mitigate these issues, we further refined the AHP images. Images were selected based on human size, with a minimum dimension of 180 pixels in height and 90 pixels in width, and standardized to a resolution of $200\times 200$. We also filtered out images with mismatched binary masks and parsing maps to prevent background bleed into human images. To address occlusions in parsing maps, we added an additional segmentation category for accessories and discarded images with disconnected parsing map parts. Our final dataset consists of about 23K images, each paired with its corresponding parsing map and OpenPose keypoints, ensuring higher consistency and quality for training.
Figure 6: Super-Resolution: lowering the impact of the fine-tuned LDM copy through $\beta$, higher quality may be achieved. When $\beta$ is too low, humans are no longer generated (column 4).
4.2 Comparison with state-of-the-art
Qualitative/quantitative analysis: Figure 5 contains our results in comparison with an unconditional GAN-based SOTA method (StyleGAN-Human [12]) as well as a diffusion-based SOTA method that has been trained using similar 2D pose conditioning as ours (ControlNet [52]) and high-resolution images. In the former, humans are generated at 1024x1024, the background is subsequently removed using rembg [2] and humans are then pasted on top of the context image, at random. In the latter, we use Stable Diffusion for inpainting where the rectangular masks are passed to the inpainting model as the ControlNet model does not provide any information about the location of the body. Our method generates more natural-looking, environment-specific humans the clothing semantics and the overall color harmonization are considered. Additionally, our quantitative analysis in Table 2 shows how our method outperforms ControlNet.
In Figure 6, we gradually decrease the impact of our fine-tuned ControlNet model through $\beta$. When $\beta$ equals 00, our system is equivalent to the vanilla Stable Diffusion inpainting model, using a binary mask. It is observed that, in this case, the output does not contain people in the desired regions. This phenomenon occurs because of a large semantic gap that exists between the complex human body and the binary inpainting mask. This experiment is to show how current, diffusion-based, state-of-the-art inpainting models, do not have the capacity to inpaint humans into existing photographs without additional guidance (HPM in our case). Numerical results are included in Table 2.
Figure 7: Pose is fixed to show how our method generates plausible clothing independent of the pose and depending on the context.
Figure 8: Results: ESP outputs a wide range of people in clothing suitable for diverse contexts. More results are included in the appendix.
Perceptual evaluation: To evaluate the appropriateness of our generated humans versus state-of-the-art, we use BLIP-2 [27] to caption a set of context images and group them by the environment type. We then select 20 prominent environment groups and randomly select one sample for each. We compete against the current state-of-the-art, i.e. a pose-conditioned LDM from [52]. 16 people were attending the short study and the results show that our method was preferred 61,875% of the time. 3.4375% of the time, the user was undecided.
4.3 Evaluation and ablation studies
Context embeddings: Figure 4 shows the blurry background reconstruction results of our (1) VAE, and (2) contextual HPM StyleGAN generator. Both predict images that are similar to the GT. StyleGAN generator’s predictions are of lower quality than the VAE. This is also noticeable in the quantitative analysis. Nevertheless, we find that these reconstructions are meaningful as they also lead to the generation of environment-inspired clothing semantics.
In Figure 3, a difference between the fixed and random context embeddings is shown in the generator. Fixing the context (rows F) generates more uniform clothing. Here, each sample is generated with a different random part of the contextual style vector, while the context embeddings are fixed. On the contrary, when the context embeddings are not fixed, our model generates more varied clothing styles (rows R). This indicates there is a link between the context and the environment, as expected.
Pose conditioning: Previously, it had been shown that pose plays an important role in the synthesis of scene-compatible humans [5]. Here we argue that the clothing semantics may also be inferred from the scene context alone. Therefore, we fix the pose in Fig. 7 to show how our method generates plausible clothing independent of the pose and depending on the image context. More samples where the original pose is used are shown in Fig. 8. For examples where the conditioning is shown, see Fig. 1.
Inpainting: Changing the impact of our ControlNet model $\beta$ may increase image quality (Figure 6). This is due to the low resolution of images that we use to train the model. When $\beta$ is lowered, the original Stable Diffusion model that was trained on high-quality images has more impact on the final result, and therefore images of higher quality are generated. However, if $\beta$ is too low, generated humans might no longer respect the input pose or they might not be generated, as discussed above.
Quantitative analysis: To determine the effectiveness of our HPM generator, we compare it to GT data with a fixed context C (denoted as F) and with a random context (R) in Table 3. The results demonstrate that our method is able to generate meaningful parsing masks and background images when the original data distribution is considered (FID, KID). It is also shown, quantitatively, that the background reconstruction errors (PSNR, MSE, LPIPS) are considerably lower when the context vector is fixed, indicating that our method generates context-aware predictions. Our reconstructed backgrounds (row 2) are similar to the VAE-reconstructed backgrounds (row 7). The VAE module provides a guiding vector that makes it possible for our HPM model to generate backgrounds that look similar to the real backgrounds (row 3). Since the HPM and the background generator share all of the parameters, we conclude that our HPM generation is context-aware. Numerical results for inpainting including the corresponding ablation of $\beta$ can be found in Table 2.
5 Discussion
Our results, as visually compelling as they are, are biased. The biases observed — such as the lack of racial diversity in certain contexts, the uniformity of skin color in scenarios like handshake interactions, the socioeconomic implications of correlating impoverished environments with darker-skinned individuals, and the gender imbalances skewed towards male representations — are a reflection of the training data’s limitations. Therefore, it is imperative to take conscious steps to mitigate these biases which includes diversifying the training data.
6 Conclusion
In this paper, we have demonstrated how full-body human generation and subsequent inpainting can be made context-aware, i.e., our generated humans are in sync with the environment when their clothing semantics and overall color harmonization are considered. Our main contribution is conditioning the image generation process on scene context to generate environment-specific people in clothing that is suitable for inpainting into a given environment. Our results demonstrate that our method produces synthetic imagery of plausible contextualized people in scenes, outperforming state-of-the-art methods on the task of contextual full-body generation.
Acknowledgement: This project has received funding from the EU Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 860768. MJB has received research gift funds from Adobe, Intel, Nvidia, Meta/Facebook, and Amazon. MJB has financial interests in Amazon, Datagen Technologies, and Meshcapade GmbH. While MJB is a consultant for Meshcapade, his research in this project was performed solely at, and funded solely by, the Max Planck Society.
Appendix
Appendix A Model architecture
Scene representation.
We extract the contextual scene representation vectors from the latent space of a variational auto-encoder (VAE). Here, the encoder consists of $5$ blocks of layers, where each block is comprised of one convolutional, batch normalization, and a leaky ReLU layer. The bottleneck layer is $512$-dimensional, and it contains our context code that is used as a conditioning in the StyleGAN generators for human parsing maps. The decoder mirrors the encoder, and its $5$ blocks are comprised of upconvolution, batch normalization, and a leaky ReLu layer. In the last block, we add an additional convolution and a tanh activation.
The architecture of the remaining modules follows the guidelines defined in Style-GAN2 [22], Stable Diffusion LDM [41], and ControlNet [52]. Code will be released.
Appendix B Training details
Scene representation.
We train the scene VAE on the AHP dataset [57] with a batch size of $64$ for approximately $500$ epochs and with a learning rate of $0.0002$. We use Adam optimizer with betas set to $0.5$ and $0.9$. The weights for the reconstruction and KL divergence loss terms are set to $\lambda_{1}=1$ and $\lambda_{2}=10$ (see Equation 1. in the main paper). After the training procedure has been completed, we freeze the VAE’s parameters and use its latent space as the context embedding to condition the contextual human parsing map generators.
The training procedure for the remaining modules follows the guidelines defined in Style-GAN2 [22], Stable Diffusion LDM [41], and ControlNet [52]. Code will be released.
Figure 9: Additional results for different contexts.
Appendix C Additional Results
Figure 9 contains more results for various environments.

### v2

@@
-Environment-Specific People
+Synthesizing Environment-Specific People in Photographs
 Abstract
-Despite significant progress in generative image synthesis and full-body generation in particular, state-of-the-art methods are either context-independent, overly reliant to text prompts, or bound to the curated training datasets, such as fashion images with monotonous backgrounds.
-Here, our goal is to generate people in clothing that is semantically appropriate for a given scene.
-To this end, we present ESP, a novel method for context-aware full-body generation, that enables photo-realistic inpainting of people into existing “in-the-wild” photographs.
-ESP is conditioned on a 2D pose and contextual cues that are extracted from the environment photograph and integrated into the generation process.
+We present ESP, a novel method for context-aware full-body generation, that enables photo-realistic synthesis and inpainting of people wearing clothing that is semantically appropriate for the scene depicted in an input photograph.
+ESP is conditioned on a 2D pose and contextual cues that are extracted from the photograph of the scene and integrated into the generation process, where the clothing is modeled explicitly with human parsing masks (HPM).
+Generated HPMs are used as tight guiding masks for inpainting, such that no changes are made to the original background.
 Our models are trained on a dataset containing a set of in-the-wild photographs of people covering a wide range of different environments.
-The method is analyzed quantitatively and qualitatively, and we show that ESP outperforms state-of-the-art on the task of contextual full-body generation.
-Figure 1: ESP generates people in semantically appropriate, environment-inspired clothing: given a 2D pose (row 1) and an environment (row 2); people are inpainted into the photo, based on a context-aware generative model that predicts contextual human parsing maps.
+The method is analyzed quantitatively and qualitatively, and we show that ESP outperforms the state-of-the-art on the task of contextual full-body generation.
+https://esp.is.tue.mpg.de/
+Keywords:
+Generative modeling Full-body synthesis Environments
+Figure 1: ESP generates people in semantically appropriate, environment-inspired clothing.
+Given a 2D pose and an environment (top), people are inpainted into the photo based on a context-aware generative model (sample 1, 2, 3).
 1 Introduction
-In recent years, great progress has been made on all aspects of image generation and editing, e.g., GAN-based methods [53, 56, 49, 28, 54, 55] allow for photo-realistic image synthesis on well-crafted datasets, and diffusion-based models [41, 52, 32] make it possible for the original image content to be seamlessly removed or replaced via inpainting.
-Due to the rapid developments in generative AI, distinguishing between AI-generated content and real content by analyzing the quality of the generated images alone is getting more difficult each day.
-In this paper, we focus on full-body generation and inpainting of environment-specific people wearing clothing that is appropriate for a given environment.
-For example, if the context of the depicted scene is a football field, our method would generate a person dressed in sportswear such as shorts, a T-shirt or a polo shirt, football boots, etc.
-However, if the context switches to snowy mountains, the inserted person would be wearing a winter jacket, ice skates, or ski attire.
+In recent years, great progress has been made on all aspects of image generation and editing.
+GAN-based methods [57, 60, 52, 30, 58, 59] allow for photo-realistic image synthesis on well-crafted datasets, and diffusion-based models [43, 55, 34] make it possible for the original image content to be seamlessly removed or replaced via inpainting.
+Rapid advancements in generative AI have resulted in tools that take text prompts and produce photo-realistic images of humans [40, 51].
+Constraining the generation requires crafting text prompts to achieve the desired result.
+In this paper, our goal is to take an image of a scene and a 2D pose and generate a realistic person, in the desired pose, wearing scene-appropriate clothing; no text prompts are required.
+For example, if the context of the depicted scene is a football field, our method generates people dressed in appropriate athletic wear such as shorts, a jersey, football helmet, etc.; see Figure 1.
+However, if the context switches to snowy mountains, the inserted person will be wearing a winter jacket, ice skates, or ski attire.
 One line of work in full-body generation is centered around using StyleGAN [12, 13] to generate full-body samples.
-Such methods are usually trained on high-quality data, such as DeepFashion [31], that is curated and hand annotated.
+Such methods are trained on high-quality data, such as DeepFashion [33], that is curated and hand annotated.
 To reduce its high dimensionality, data is first normalized and filtered.
 This is a crucial step that is contributing to the success of StyleGAN-based methods.
-While the images that are generated in this way are of high-quality, the main issue that we tackle in this paper is the semantic asyncrony between the clothing and the corresponding environment that exists in these datasets due to the unnatural data acquisition settings.
-I.e., in fashion photography, images are usually captured in a studio setting, where the subject is fully lit and the background is very simple, e.g., a white screen.
-Therefore there is minimal environmental context available.
-Also, in case of an outdoor photo-shoot, the link between the environment and the clothing may be atypical and it does not represent what is normally worn on the location well.
-The other, more recent line of work, are diffusion-based methods or, typically, large pre-trained latent diffusion models such as Stable Diffusion [41] and its conditional variations (e.g.
+While the images that are generated in this way are of high-quality, the generated clothing is typically not correlated with the scene due to the unnatural settings present in the training data.
+For example, in fashion photography, images are usually captured in a studio setting, where the subject is fully lit and the background is very simple, e.g., a white screen.
+Therefore, there is minimal environmental context available.
+The more recent line of work focuses on diffusion-based methods that generate images from text, e.g.
+large pre-trained latent diffusion models (LDM) such as Stable Diffusion [43] and its conditional variations (e.g.
 ControlNet [55]).
 The main advantage of such methods is the existence of large pre-trained image priors that are usually fine-tuned on human data including full-bodies in the wild.
-However, in order to generate sufficiently varied samples, these txt2img models require accurate textual descriptions of the desired output.
-While providing such descriptions may lead to satisfactory results, in this work we assume that there are no textual descriptions of the appropriate clothing available.
-And instead, we focus on visual information alone, investigating whether the image context can be used as a sole driving source for the generation of environment-specific people.
+However, in order to generate sufficiently varied samples, these text-to-image (T2M) models require accurate textual descriptions of the desired output.
+In contrast, here we seek an automated solution and do not assume that textual descriptions of the appropriate clothing are available.
+Instead, we focus on visual information alone, investigating whether the image context can be used as a sole driving source for the generation of environment-specific people.
 To this end, we take advantage of both StyleGAN-based and diffusion-based methods, combining their strengths to devise a purely image-based system for our task.
-As part of our pipeline, contextual cues are first extracted from a given environment photo with a VAE and then fed into a StyleGAN-based human parsing map (HPM) generator that predicts context-aware clothing semantics.
-HPM training data is filtered and normalized to include only simple poses, making it suitable for training a StyleGAN model.
-Next, a HPM translation module is fine-tuned using the same dataset and in turn utilized to provide additional guidance to the inpainting process of a large pre-trained Stable Diffusion model, for seamless inpainting of photo-realistic and environment-specific people into the original photographs.
-Our method is evaluated quantitatively and qualitatively, with a focus on contextual appropriateness.
-The results show that our method generates environment-specific people that match the context of the input photograph well, considering the clothing semantics (see Figure 1).
-Furthermore, our approach outperforms state of the art on the task of contextual full-body generation.
-Our contributions are as follows:
+As part of our pipeline, contextual cues are first extracted from a given environment photo with a VAE and, then, fed into a StyleGAN-based human parsing map (HPM) generator that predicts both clothing semantics and the original environment.
+The key insight is that, when combined with a random vector, the extracted contextual cues form our contextual-style vectors that allow for a stochastic generation of a large variety of context-appropriate clothing.
+These vectors are connected back to the original environment, as they jointly drive the generation of both the HPMs and the corresponding environment, making the HPM generation context-aware.
+To disentangle the pose from the context, the generator is further conditioned on 2D pose renderings and the network’s computational graph is extended through an image-to-image (I2I) translation module (similar to pix2pix [20]).
+Finally, leveraging a large pre-trained T2M LDM for its state-of-the-art image synthesis and inpainting capacities [43], generated HPMs are used as input conditioning and guidance during the inpainting and super-resolution process.
+Our method is evaluated quantitatively and qualitatively, demonstrating that our approach outperforms the state of the art on the task of contextual full-body generation based on the following contributions:
+  • $\circ$
     A StyleGAN-based method for stochastic generation of context-aware HPMs that are generated using our newly introduced contextual style vectors;
-  • a dynamic pix2pix-based extension of the computational graph that is fine-tuned in parallel with the StyleGAN generator which enables pose-guided imgage-to-image translation;
-  • a system that leverages a pre-trained text-to-image latent diffusion model (Stable Diffusion), allowing for super-resolution imaging and photorealistic inpainting of environment-specific people into existing photographs.
-Figure 2: System Overview: (I) an input context image is encoded into the latent space of a VAE, giving a contextual style vector; (II) the latter is then fed, alongside a random vector, into a pose-conditioned StyleGAN-II HPM generator; (III) an HPM translation module is used to achieve fine-grained control over the generated clothing during the inpainting process (using the generic Stable Diffusion LDM).
+  • $\circ$
+    An extension of the computational graph that is trained end-to-end with the StyleGAN generator which enables pose-guided I2I translation;
+  • $\circ$
+    A system that leverages a large pre-trained T2M latent diffusion model (Stable Diffusion), allowing for super-resolution imaging and inpainting of the environment-specific people into the existing photographs.
 2 Related work
-From their advent, Generative Adversarial Networks (GAN) [13] have proven to be a very effective tool for synthesizing realistic images.
+Generative Adversarial Networks (GAN) [14] have proven to be a very effective tool for synthesizing realistic images.
 StyleGAN-based methods [24, 22, 23, 5] have been used to generate photo-realistic imagery of a variety of objects and scenes [53, 54], including humans [28, 47].
-More recently, large latent diffusion models have been in the spotlight [37] due to their unparalled image generation capabilities, conditioned on text (txt2img)  [41, 38, 39, 1, 42] and additional input controls  [52].
+More recently, large latent diffusion models have been in the spotlight [39] due to their unparalled image generation capabilities, conditioned on text (T2M)  [43, 41, 40, 2, 44] and other controls  [55].
 Full-body generation:
-StyleGAN-based state-of-the-art for unconditional full-body generation include InsetGAN [11] and StyleGAN-Human [12].
-Both methods use well prepared datasets that are either proprietary or contain fashion imagery.
+StyleGAN-based state-of-the-art methods for unconditional full-body generation include InsetGAN [12] and StyleGAN-Human [13].
+Both methods use well-prepared datasets that are either proprietary or contain fashion imagery.
 InsetGAN composes multiple GANs to handle the complex task of generating high resolution full body human images.
@@
 The background in this data is often white or plain, or the link between the background and the clothing is atypical.
-We use a small dataset of in-the-wild imagery [57] for training, where people are clothed synchronously with the environment, e.g.
-football gear is worn on a football field and warm winter attire is worn in the snow.
+We use a small dataset of in-the-wild imagery [61] for training, where people are clothed according to the environment, e.g., football gear is worn on a football field and warm winter attire in the snow.
 Pose and/or appearance information are used to control the process of conditional full-body generation [4, 35, 36, 49, 46].
-In a two step process, Lewis et al.
+In a two-step process, Lewis et al.
 [28] first train a dense-pose conditioned StyleGAN2 model with a clothing segmentation branch, after which a new image is synthesized using layered latent space interpolation to transfer a garment to a target body.
@@
 Our twist considering this line of work is to tackle the problem from a novel perspective, by incorporating image context as a conditioning through the mapping network.
-In this way, we generate environemnt-specific people that may be inpainted into the existing photographs, while respecting the scene context in terms of the clothing semantics.
+In this way, we generate environment-specific people that may be inpainted into the existing photographs, while respecting the scene context in terms of the clothing semantics.
 Inpainting:
@@
 However, this approach slows down the image generation process of a DDPM.
-More recently, similar ideas in the realm of latent diffusion models have been explored using denoising diffusion implicit models (DDIM) [48] where the denoising steps are skipped for faster inference.
+More recently, similar ideas in the realm of latent diffusion models have been explored using denoising diffusion implicit models (DDIM) [50], where the denoising steps are skipped for faster inference.
 Most recent image generators that have the inpainting capacity use this principle and we take advantage of these models  [43].
@@
 Brooks et al.
-[5] generate both the scene and a pose-compatible human that matches the scene context at 128x128, where the context embeddings are calculated from a 2D pose vector and a noise vector.
+[6] generate both the scene and a pose-compatible human that matches the scene context at $128\times 128$, where the context embeddings are calculated from a 2D pose vector and a noise vector.
 In their work, the focus is on generating both compatible scenes and humans inside, conditioned on 2D pose embeddings.
@@
 Kulal et al.
-[25] filter a large-scale video dataset (2.4 million clips) at 256x256 and use it to train a multi-faceted LDM, generating pose-compatible people conditioned on a photo of a reference person and a masked scene.
+[27] filter a large-scale video dataset (2.4 million clips) at $256\times 256$ and use it to train a multi-faceted LDM, generating pose-compatible people conditioned on a photo of a reference person and a masked scene.
 The primary focus of this work is on synthesizing plausible poses given a scene context.
@@
 Additionally, we only use a small dataset containing approx.
-23K images at 200x200, and our results are generated at 512x512.
+23K images at $200\times 200$, and our results are generated at $512\times 512$.
 Image quality of our method is on par with state-of-the-art image generators including ControlNet [55] and Stable Diffusion [43] that many other recent high-quality full-body generation methods build on top of  [32].
+Figure 2: System overview: (I) an input context image is encoded into the latent space of a VAE, giving context embeddings; (II) the latter is then fed, alongside a random vector, creating a contextual style vector, into a pose-conditioned StyleGAN-II HPM generator; (III) generated HPMs are used as input for pretrained Stable Diffusion/CN modules to achieve fine-grained control over the generated clothing during inpainting.
 3 Method
-Our system is shown in Figure 2.
+Our method is shown in Figure 2.
 First, context embeddings are extracted from image $\mathcal{I}$ that contains the environmental context $\mathcal{C}$.
-Then, using the extracted embeddings, small human parsing maps $\mathcal{M}$ are generated conditioned on a 2D skeleton rendering $\mathcal{S}$ (32x64).
-Next, generated HPMs are then used, with the help of our HPM translation module, as an additional guiding signal for the seamless inpainting of environment-specific people into the existing photographs.
-In this final step, we take advantage of a large pre-trained LDM (Stable Diffusion for inpainting), allowing for the generation of images at very high resolution (512x512).
+Then, using the extracted embeddings, human parsing maps (HPMs) $\mathcal{M}$ are generated conditioned on a 2D skeleton rendering $\mathcal{S}$ ($32\times 64$).
+The generated HPMs are then used, with the help of our HPM translation module, as an additional guiding signal for inpainting of environment-specific people into the existing photographs.
+In this final step, we take advantage of a large pre-trained LDM (Stable Diffusion for inpainting), allowing for the generation of images at high resolution ($512\times 512$).
 3.1 Context embeddings
@@
 We dilate the mask with morphological operations and invert it to get the soft background mask $M_{b}$.
-To remove the foreground human, we apply a pixel-wise mask on the input image, producing the masked image $\mathcal{I}=I_{c}^{h}\odot M_{b}$ Variational auto-encoders (VAE) [24] are good at representing the overall global information present in an image [15], so we choose the compact and normalized latent space of a VAE to reflect the global structure, illumination and color tone of the context image $\mathcal{I}$.
-We train the VAE with $i_{c}$ where $i_{c}$ is a down-sampled $64\times 64$ version of $\mathcal{I}$ with the loss:
+To remove the foreground human, we apply a pixel-wise mask on the input image, producing the masked image $\mathcal{I}=I_{c}^{h}\odot M_{b}$
+We choose the compact and normalized latent space of a Variational Auto-Encoder (VAE) [26, 16] to reflect the global structure, illumination and color tone of the context image $\mathcal{I}$.
+We train the VAE with $i_{c}$, where $i_{c}$ is a down-sampled $64\times 64$ version of $\mathcal{I}$ with the loss:
 $\begin{split}L_{V}=\lambda_{1}\ ||i_{c}-\tilde{i}_{c}||_{1}+\lambda_{2}D_{KL}\ ({\scriptstyle\mathcal{N}(\mu_{c},\sigma^{2}_{c})}\ ||\ {\scriptstyle\mathcal{N}(0,1)}),\end{split}$ (1)
-where $\tilde{i}_{c}$ is the reconstructed image from the VAE; $\lambda$’s are the corresponding weights for the different losses; $D_{KL}$ is the KL divergence loss [24]; $\mathcal{N}(x,y)$ is the normal distribution with mean $x$ and variance $y$; and $\mu_{c}$, $\sigma^{2}_{c}$ are the mean and variance of the context image.
-Once trained, the latent vector $\mathcal{C}$ sampled from the VAE for corresponding context image $\mathcal{I}$ is used for conditioning in the following steps.
-Note that the reconstruction loss term is applied only to the regions where humans are present and as such we do not require $\mathcal{I}$ to be computed from $I_{c}^{h}$ at inference time.
-Additionally, $\tilde{i}_{c}$ is subsequently passed through the VAE to remove the small artifacts that may appear in the regions that people typically occupy in our training data.
+where $\tilde{i}_{c}$ is the reconstructed image from the VAE; $\lambda$s are the corresponding weights for the different losses; $D_{KL}$ is the KL divergence loss [26]; $\mathcal{N}(x,y)$ is the normal distribution with mean $x$ and variance $y$; and $\mu_{c}$, $\sigma^{2}_{c}$ are the mean and variance of the context image.
+Once trained, the latent vector $\mathcal{C}$ sampled from the VAE for the corresponding context image $\mathcal{I}$ is used for conditioning in the following steps.
+Note that the reconstruction loss term is applied only to the regions where humans are present, and, as such, we do not require $\mathcal{I}$ to be computed from $I_{c}^{h}$ at inference time.
 3.2 Contextualized generation of HPMs
@@
 [20], to inject the pose conditioning into the generator, we concatenate the input of different style blocks of the generator with the output of the corresponding layers of a pose encoder.
-The input to the encoder is an RGB image $\mathcal{S}$ and its layers are designed so that the number of the feature maps and their dimensions match the StyleGAN2 generator (32x64).
+The input to the encoder is an RGB image $\mathcal{S}$ and its layers are designed so that the number of the feature maps and their dimensions match the StyleGAN2 generator ($32\times 64$).
 The generator and the pose encoder are trained end-to-end.
 The HPM generator jointly outputs a HPM $\tilde{\mathcal{M}}$ (3 channels) and a blurry background image ($3$ channels).
-Note that the blurry background reconstruction from the VAE that is used as the GT image for the generator is rescaled from 64x64 to 32x64 to match the HPM image dimensions.
-Similar to [11], we observe that GANs do not handle large areas of uniform color well and therefore we use the tall HPM images where the number of the background pixels has been reduced, rather than square images with many background pixels.
+Note that the blurry background reconstruction from the VAE that is used as the GT image for the generator is rescaled from $64\times 64$ to $32\times 64$ to match the HPM image dimensions.
+Similar to [12], we observe that GANs do not handle large areas of uniform color well and, therefore, we use the tall HPM images, where the number of the background pixels has been reduced, rather than square images with many background pixels.
 We are able to do so because our dataset is filtered and it only contains the poses where the width of the person is equal to at most one half of their height.
@@
 In summary, the generator loss for the semantic maps is defined as:
-$\begin{split}L_{S}^{G}=L_{adv}^{D_{\mathcal{M}}}+L_{adv}^{D_{B}}+L_{adv}^{D_{MB}}\\ +\lambda_{3}L_{1}+\lambda_{4}L_{R}+\lambda_{5}L_{perc}\end{split}$ (2)
-where $L_{adv}^{D_{\mathcal{M}}}$ and $L_{adv}^{D_{B}}$ represent the adversarial losses that come from the HPM and the blurry background image discriminators respectively; $L_{adv}^{D_{MB}}$ is the loss from an additional discriminator that evaluates the composite of the blurry background and the HPM; $L_{1}$ is the $l_{1}$ loss applied between the generated blurry background and the corresponding generated background $\tilde{i}_{c}$ from the VAE; $L_{perc}$ is a standard perceptual loss commonly used in the literature for generating more detailed images; $L_{R}$ represents the path regularisation [22] that is applied to the generator; $\lambda_{3}$, $\lambda_{4}$, and $\lambda_{5}$ are the weighting factors for the specific losses.
+$\begin{split}L_{S}^{G}=L_{adv}^{D_{\mathcal{M}}}+L_{adv}^{D_{B}}+L_{adv}^{D_{MB}}+\lambda_{3}L_{1}+\lambda_{4}L_{R}+\lambda_{5}L_{perc},\end{split}$ (2)
+where $L_{adv}^{D_{\mathcal{M}}}$ and $L_{adv}^{D_{B}}$ represent the adversarial losses that come from the HPM and the blurry background image discriminators respectively; $L_{adv}^{D_{MB}}$ is the loss from an additional discriminator that evaluates the composite of the blurry background and the HPM; $L_{1}$ is the $l_{1}$ loss applied between the generated blurry background and the corresponding generated background $\tilde{i}_{c}$ from the VAE; $L_{perc}$ is a standard perceptual loss commonly used in the literature for generating more detailed images [56]; $L_{R}$ represents the path regularisation [25] that is applied to the generator; $\lambda_{3}$, $\lambda_{4}$, and $\lambda_{5}$ are the weighting factors for the specific losses.
 3.3 HPM translation module
 In the remainder of the pipeline, we use diffusion models [18].
-In particular, we leverage the large pre-trained latent diffusion model that is called Stable Diffusion [41] as our image prior.
+In particular, we leverage the large pre-trained latent diffusion model Stable Diffusion [43] as our image prior.
 Diffusion models are well-known for their capacity to generate highly detailed images.
-They do so gradually through the denoising diffusion process in $T$ timesteps.
-During training, Gaussian noise $\mathcal{N}(\mathbf{0},\mathbf{I})$ gets added to the original sample $\mathbf{x}_{0}$, step by step, and a model similar to UNet [18] is trained to predict the noise $\epsilon$ residuals, at timestep $t$.
-For latent diffusion models where the denoising process is performed in the latent space rather than in pixel space, $\mathbf{z}_{t}$ marks the corresponding latent representation of a noisy sample $\mathbf{x}_{t}$ at timestep $t$.
+They do this gradually, through the denoising diffusion process in $T$ timesteps.
+During training, Gaussian noise $\mathcal{N}(\mathbf{0},\mathbf{I})$ gets added to the original sample $\mathbf{x}_{0}$, step by step, and a UNet-like model [19] is trained to predict the noise $\epsilon$ residuals, at timestep $t$.
+For latent diffusion models (LDM), where the denoising process is performed in the latent space rather than in pixel space, $\mathbf{z}_{t}$ marks the corresponding latent representation of a noisy sample $\mathbf{x}_{t}$ at timestep $t$.
 The LDM objective is then formulated as follows:
@@
 Given the large success of LDMs, new mechanisms have appeared on how to control their image generation process further.
-In this paper we use the approach called ControlNet from [52].
+In this paper, we use the ControlNet [55] as a backbone.
 The main idea behind ControlNet is to create a trainable copy of a pre-trained LDM, attach it to the LDM and in turn train it without modifying the original model (zero-convolutions).
@@
 The copy is additionally conditioned on images $\mathbf{c}_{\text{f}}$ that are encoded into its latent space by a small learnt neural network and the feature maps of the encoder of the LDM copy are added to the feature maps of the original LDM so as to transfer the original conditioning signal.
-Including the text conditioning $\mathbf{c}_{t}$, objective fuction for a conditional LDM is now equal to:
+Including the text conditioning $\mathbf{c}_{t}$, the objective function for a conditional LDM is now equal to:
 $\vspace{-3pt}\begin{split}L^{C}=\mathbb{E}_{\mathbf{z}_{0},\mathbf{t},\mathbf{c}_{t},\mathbf{c}_{\text{f}},\epsilon\sim\mathcal{N}(0,1)}\Big{[}\|\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},\mathbf{t},\mathbf{c}_{t},\mathbf{c}_{\text{f}}))\|_{2}^{2}\Big{]}\,.\end{split}$ (4)
 We attach a ControlNet module to the Stable Diffusion model and train it on approx.
-23K images where the input conditioning are RGB HPM images while the output are the corresponding images of real people in the wild.
-This model is later used as a translator for our custom low resolution RGB HPM images that are padded and then upscaled from 32x64 to 512x512 during training.
-The output is upscaled from 200x200 to 512x512 during training.
+$23K$ images, where the input conditioning are RGB HPM images and the output are the corresponding images of real people in the wild.
+During training, we use a generic text prompt that is dropped $50\%$ of the time.
+This model is later used as a translator for our custom low-resolution RGB HPM images that are padded and then upscaled from $32\times 64$ to $512\times 512$ during training.
+The output is upscaled from $200\times 200$ to $512\times 512$ during training.
+Figure 3: Analysis of the StyleGAN generator: A.
+Contextual HPM Generation: Fixing the context (rows F) generates more uniform clothing than varying the context embeddings (rows R).
+This indicates that there is a link between the context and the environment that has been learned by our context-aware HPM generator.
+B.
+Background Reconstruction: GT/VAE reconstructions are shown in rows 1/2; fixing the context vector in the StyleGAN generator gives matching results (row F) with GT/VAE, while varying the context embeddings leads to random predictions (row R).
 3.4 Inpainting and super-resolution
-Human body is highly complex due to its high flexibility that stems from the many degrees of freedom that it has.
-Also, one body may be clothed in many different types of outfits.
+The human body is highly complex due to its large range of poses and also appearances, as one body may be clothed in many different types of outfits.
 For this reason, there exists a large semantic gap between the body and its binary segmentation mask.
@@
 In this way, we are able to transfer the background from the original image into the resulting image using the inpainting model while predicting the HPM-compatible humans at the same time.
-Additionally, by using a weighting scheme where the impact of the HPM translator is reduced, it is possible to generate images of very high quality where the quality is similar to that of the Stable Diffusion-generated imagery.
-This is done by reducing the weight $\beta$ (controlNet strength) that is used for the multiplicaton of the ControlNet’s feature maps prior to their addition with the feature maps of the inpainting model.
-However, $\beta$ should not be reduced drastically as the benefits of the guidance signal will become undone.
-We find that setting $\beta$ to $0.5$ works well.
-Figure 3: Fixing the context (rows F) generates more uniform clothing than varying the context embeddings (rows R).
-This indicates there is a link between the context and the environment that has been learned by our context-aware HPM generator.
-Figure 4: Background Reconstruction: VAE reconstructions are shown in rows 1 (GT) and 2 (VAE); fixing the context vector in the StyleGAN generator gives matching results (row F) with GT/VAE, while varying the context leads to random predictions (row R).
-Figure 5: SOTA Comparison: the environment where the humans are to be inpainted is shown in the first row; results with StyleGAN-Human [12] are in row 2; ControlNet [52] results are in row 3.
-Our method generates more appropriate, environment-specific, people.
-Output $\psi=0.5$ $\psi=0.6$ $\psi=0.7$ $\psi=0.8$ $\psi=0.9$ $\psi=1$
-HPM 15.73 14.59 13.53 12.64 11.96 11.49
-Context 112.45 86.33 67.55 54.77 46.29 41.08
-Table 1: Truncation Trick: the table shows FIDnormal-$\downarrow$ scores for varying $\psi$ across 10K randomly selected contexts.
-The optimal value of $\psi$ is equal to $1.0$ for both the HPM and the Context generator.
-Method $\beta=0$ $\beta=0.25$ $\beta=0.5$ $\beta=0.75$ $\beta=1$
-Baseline - - 78.7 86.7 90.0
-Ours 42.3 22.6 21.0 19.8 20.2
-Table 2: ControlNet Strength: the table shows FIDnormal-$\downarrow$ scores for varying $\beta$ across 10K images. Our method outperforms SOTA.
-4 Experiments
-4.1 Data
-Our method utilizes a training dataset comprising context images, images with humans, skeletal pose information, and machine-annotated human body parsing or semantic segmentation maps.
-For training with the AHP dataset [57], we extract context images by using segmentation masks to remove people from the images.
-Skeleton images with pose data are generated using OpenPose [7, 6] on these human-included context images.
-During inference, only a context image and pose information are required to generate a full-body person image with context-aware appearance.
+The impact of the HPM translator can be adjusted with a weight $\beta$ (controlNet strength) that is used for the multiplication of the ControlNet’s feature maps prior to their addition with the feature maps of the inpainting model.
+A low weight leads to a reduced guidance while producing higher quality imagery.
+We found that setting $\beta$ to $0.5$ leads to satisfactory results.
+Table 1: Quantitative analysis: C denotes whether the context vector is fixed (F) or random (R).
+Fixing the context vector leads to predictions that are closer to GT, especially in background reconstruction.
+The results are calculated over 10K contexts.
+The analysis confirms that our method is context-aware (PSNR, MSE, LPIPS) and generates meaningful HPMs and backgrounds (FID, KID).
+The performance of the VAE module used to generate the blurry contexts (rows 2 and 5) is shown in row 7.
 C Evaluation PSNR $\uparrow$ MSE $\downarrow$ LPIPS $\downarrow$ FID $\downarrow$ KID $\downarrow$
@@
 F VAE Recon. 27.05 0.0022 0.051 7.0 0.0029
-Table 3: Quantitative Analysis: the results are calculated over 10K contexts.
-C denotes whether the context vector is fixed (F) or random (R).
-Overall, fixing the context vector leads to the predictions that are closer to the ground truth, especially in background reconstruction.
-Therefore, the analysis confirms that our method is context-aware (PSNR, MSE, LPIPS) and that it generates meaningful parsing masks and background contexts (FID, KID).
-The performance of the standalone VAE module used to generate the blurry contexts (rows 2 and 5) is shown in the last row.
+4 Experiments
+Figure 4: Comparison against state of the art: we compare (A) StyleGAN-Human, (B) TextHuman, (C) ControlNet (OpenPose) & SD-2.1-Inpainting, with (D) Our full method.
+(A) is unconditional, while (B) is conditioned on a dense pose.
+Both do not respect the context - considering clothing semantics, and lighting.
+Parts of the body may be missing (see (B), col 2).
+While (C) is conditioned on a pose, the pose is not always respected.
+It changes the original scene (e.g.
+by insertion of new objects; see col 3) and it does not always generate single humans (see cols 1, 5).
+In contrast, Ours (D) does not change the scene, it generates single humans that are posed correctly, and overall contextual alignment, when clothing semantics and lighting are considered, is higher than in (A), (B), and (C).
+Note: (B), (C), and (D) have the same input pose.
+4.1 Data
+Our method utilizes the AHP dataset [61] comprising context images, images with humans, machine-annotated 2D pose information, and HPMs.
+For training, we extract context images by using segmentation masks to remove people from the images.
+Skeleton images with pose data are generated using OpenPose [7, 8].
+During inference, only a context image and the pose information are required.
 The AHP dataset is compiled from large-scale instance segmentation and detection datasets, including COCO [31], VOC [11], SBD [17], and Objects365 [48].
-It underwent filtering and post-processing to yield 56,599 images with parsing maps and binary human masks, which are integral to our method.
-However, the raw AHP dataset’s varied resolution and quality, along with inconsistent human pixel percentages, lead to suboptimal results.
+It underwent filtering and post-processing to yield 56K images with parsing maps and binary human masks, which are integral to our method.
+However, the raw AHP dataset’s varied resolution and quality, along with inconsistent human pixel percentages, lead to sub-optimal results.
 Additionally, the provided parsing maps sometimes have disconnections and holes.
@@
 Images were selected based on human size, with a minimum dimension of $180$ pixels in height and 90 pixels in width, and standardized to a resolution of $200\times 200$.
-We also filtered out images with mismatched binary masks and parsing maps to prevent background bleed into human images.
-To address occlusions in parsing maps, we added an additional segmentation category for accessories and discarded images with disconnected parsing map parts.
-Our final dataset consists of about 23K images, each paired with its corresponding parsing map and OpenPose keypoints, ensuring higher consistency and quality for training.
-Figure 6: Super-Resolution: lowering the impact of the fine-tuned LDM copy through $\beta$, higher quality may be achieved. When $\beta$ is too low, humans are no longer generated (column 4).
-4.2 Comparison with state-of-the-art
-Qualitative/quantitative analysis: Figure 5 contains our results in comparison with an unconditional GAN-based SOTA method (StyleGAN-Human [12]) as well as a diffusion-based SOTA method that has been trained using similar 2D pose conditioning as ours (ControlNet [52]) and high-resolution images.
-In the former, humans are generated at 1024x1024, the background is subsequently removed using rembg [2] and humans are then pasted on top of the context image, at random.
-In the latter, we use Stable Diffusion for inpainting where the rectangular masks are passed to the inpainting model as the ControlNet model does not provide any information about the location of the body.
-Our method generates more natural-looking, environment-specific humans the clothing semantics and the overall color harmonization are considered.
-Additionally, our quantitative analysis in Table 2 shows how our method outperforms ControlNet.
-In Figure 6, we gradually decrease the impact of our fine-tuned ControlNet model through $\beta$.
-When $\beta$ equals 00, our system is equivalent to the vanilla Stable Diffusion inpainting model, using a binary mask.
-It is observed that, in this case, the output does not contain people in the desired regions.
-This phenomenon occurs because of a large semantic gap that exists between the complex human body and the binary inpainting mask.
-This experiment is to show how current, diffusion-based, state-of-the-art inpainting models, do not have the capacity to inpaint humans into existing photographs without additional guidance (HPM in our case).
-Numerical results are included in Table 2.
-Figure 7: Pose is fixed to show how our method generates plausible clothing independent of the pose and depending on the context.
-Figure 8: Results: ESP outputs a wide range of people in clothing suitable for diverse contexts. More results are included in the appendix.
+We also filtered out images with mismatched binary masks and parsing maps to prevent background bleeds into human images.
+To address occlusions in parsing maps, we added a segmentation category for accessories and discarded images with disconnected parsing map parts.
+Our final dataset consists of $\sim 23$K images, each paired with its corresponding parsing map and OpenPose keypoints, ensuring higher consistency and quality for training.
+[…]
+Output $\psi=0.5$ $\psi=0.6$ $\psi=0.7$ $\psi=0.8$ $\psi=0.9$ $\psi=1$
+HPM 15.73 14.59 13.53 12.64 11.96 11.49
+Context 112.45 86.33 67.55 54.77 46.29 41.08
+[…]
+Method $\beta=0$ $\beta=0.25$ $\beta=0.5$ $\beta=0.75$ $\beta=1$
+Baseline - - 78.7 86.7 90.0
+Ours 42.3 22.6 21.0 19.8 20.2
+4.2 Comparison to state of the art
+Qualitative/quantitative analysis: Figure 4 contains a comparison with state-of-the-art full-body generation methods including (1) an unconditional GAN-based method (StyleGAN-Human [13]), (2) dense pose conditioned GAN-based method (TextHuman [21]) as well as (3) a 2D pose conditioned diffusion-based method (ControlNet [55]) using the same inpainting module [43] as in Ours.
+In (1), humans are generated at $1024\times 1024$, the background is subsequently removed using rembg [3] and humans are then pasted on top of the context image, at random.
+Next, in (2), humans are generated at $512\times 512$ from their randomized HPM predictions, and the same procedure is followed as in (1).
+In (3), rectangular masks are passed to the inpainting model as the 2D pose-conditioned ControlNet model does not provide information about the location of the body/clothing.
+Humans are inpainted at $512\times 512$.
+Similarly to (3), our method inpaints humans using a diffusion model, however, we use GAN-generated HPM masks for inpainting and additional guidance.
+Our method generates more natural-looking, environment-specific humans when the clothing semantics and the overall color harmonization are considered.
+Table 3 shows quantitative analysis against the 2D pose-conditioned ControlNet baseline.
+In Figure 5 (A), we gradually decrease the impact of our fine-tuned ControlNet model through $\beta$.
+When $\beta=0$, our system is equivalent to the vanilla Stable Diffusion inpainting model, using a binary mask (from the predicted HPM).
+It fails to inpaint humans into the photographs without additional guidance.
+Figure 5: A.
+Super-Resolution: higher image quality may be achieved by lowering the ControlNet strength via $\beta$.
+However, when $\beta$ is too low, humans are no longer generated ($\beta=0$ shows the vanilla SD inpainting result without additional HPM guidance).
+B.
+Fixed pose: generated clothing changes depending on the context.
+Figure 6: ESP outputs full-body humans wearing environment-inspired clothing that is semantically suitable for diverse contexts.
+We show a wide range of different scenes including indoors and outdoors, with varying image resolutions (from $200\times 200$ to $512\times 512$), target pose, and lighting conditions.
+The produced results depend on the quality of the context image and on the capabilities of the generic Stable Diffusion 2.1 inpainting model that is used in the final step of our system.
+Please zoom in for details.
+Figure 7: Out-of-domain generalization: our results on top of synthetic backgrounds.
 Perceptual evaluation: To evaluate the appropriateness of our generated humans versus state-of-the-art, we use BLIP-2 [29] to caption a set of context images and group them by the environment type.
 We then select 20 prominent environment groups and randomly select one sample for each.
-We compete against the current state-of-the-art, i.e.
+We compare against the current state-of-the-art, i.e.
 a pose-conditioned LDM from [55].
-16 people were attending the short study and the results show that our method was preferred 61,875% of the time.
-3.4375% of the time, the user was undecided.
+16 people attended the short study and the results show that our method was preferred 62% of the time.
+3% of the time, the user was undecided.
 4.3 Evaluation and ablation studies
-Context embeddings: Figure 4 shows the blurry background reconstruction results of our (1) VAE, and (2) contextual HPM StyleGAN generator.
-Both predict images that are similar to the GT.
-StyleGAN generator’s predictions are of lower quality than the VAE.
-This is also noticeable in the quantitative analysis.
+Quantitative analysis: To determine the effectiveness of our HPM generator, we compare it to GT data with a fixed context C (denoted as F) and with a random context (R) in Table 1.
+The results demonstrate that our method generates meaningful parsing masks and background images when the original data distribution is considered (FID, KID).
+It is also shown that the background reconstruction errors (PSNR, MSE, LPIPS) are considerably lower when the context vector is fixed because our method generates context-aware predictions.
+The reconstructed backgrounds (row 2) are similar to the VAE predictions (row 7).
+The VAE module provides embeddings that allow for the HPM model to generate backgrounds that look similar to the real ones (row 3).
+Finally, Table 2 shows the ablation of the truncation trick $\psi$ parameter from  [25].
+Context embeddings: Figure 3 (B) shows the blurry background reconstructions of our (1) VAE, and (2) contextual HPM StyleGAN generator.
+StyleGAN’s predictions are of lower quality than the VAE.
 Nevertheless, we find that these reconstructions are meaningful as they also lead to the generation of environment-inspired clothing semantics.
-In Figure 3, a difference between the fixed and random context embeddings is shown in the generator.
+In Figure 3 (A), the difference between the fixed and random context embeddings is shown in the generator.
 Fixing the context (rows F) generates more uniform clothing.
@@
 On the contrary, when the context embeddings are not fixed, our model generates more varied clothing styles (rows R).
-This indicates there is a link between the context and the environment, as expected.
-Pose conditioning: Previously, it had been shown that pose plays an important role in the synthesis of scene-compatible humans [5].
-Here we argue that the clothing semantics may also be inferred from the scene context alone.
-Therefore, we fix the pose in Fig.
-7 to show how our method generates plausible clothing independent of the pose and depending on the image context.
-More samples where the original pose is used are shown in Fig.
-8.
-For examples where the conditioning is shown, see Fig.
-1.
-Inpainting: Changing the impact of our ControlNet model $\beta$ may increase image quality (Figure 6).
+This indicates there is a link between the context and the environment.
+Inpainting: Changing the impact of our ControlNet model $\beta$ may increase image quality (Figure 5 (A)).
 This is due to the low resolution of images that we use to train the model.
 When $\beta$ is lowered, the original Stable Diffusion model that was trained on high-quality images has more impact on the final result, and therefore images of higher quality are generated.
-However, if $\beta$ is too low, generated humans might no longer respect the input pose or they might not be generated, as discussed above.
-Quantitative analysis: To determine the effectiveness of our HPM generator, we compare it to GT data with a fixed context C (denoted as F) and with a random context (R) in Table 3.
-The results demonstrate that our method is able to generate meaningful parsing masks and background images when the original data distribution is considered (FID, KID).
-It is also shown, quantitatively, that the background reconstruction errors (PSNR, MSE, LPIPS) are considerably lower when the context vector is fixed, indicating that our method generates context-aware predictions.
-Our reconstructed backgrounds (row 2) are similar to the VAE-reconstructed backgrounds (row 7).
-The VAE module provides a guiding vector that makes it possible for our HPM model to generate backgrounds that look similar to the real backgrounds (row 3).
-Since the HPM and the background generator share all of the parameters, we conclude that our HPM generation is context-aware.
-Numerical results for inpainting including the corresponding ablation of $\beta$ can be found in Table 2.
-5 Discussion
-Our results, as visually compelling as they are, are biased.
-The biases observed — such as the lack of racial diversity in certain contexts, the uniformity of skin color in scenarios like handshake interactions, the socioeconomic implications of correlating impoverished environments with darker-skinned individuals, and the gender imbalances skewed towards male representations — are a reflection of the training data’s limitations.
-Therefore, it is imperative to take conscious steps to mitigate these biases which includes diversifying the training data.
+However, if $\beta$ is too low, generated humans might no longer respect the input pose or they might not be generated, as discussed above in Sec.
+4.2.
+For numerical analysis, see Table 3.
+Pose conditioning: Previously, it had been shown that pose plays an important role in the synthesis of scene-compatible humans [6].
+Here, we argue that the clothing semantics may also be inferred from the scene context alone.
+We fix the pose in Fig.
+5(B) and Fig.
+7 to show how our method generates plausible clothing.
+More examples with the original pose are shown in Fig.
+1 and Fig.
+6.
+Limitations: Our method generates people in simple poses only (see Sec.
+4), and, if the conditioning pose is not plausible, the method might fail to generate a valid HPM.
+Also, the quality of our results depends on the capabilities of the generic inpainting model and may be further enhanced using an upscaler (Fig.
+8).
+Potential negative impact: Our system is biased as there is a lack of racial diversity in specific contexts, and the gender imbalances skewed toward male representations are a reflection of the training data’s limitations.
+Therefore, it is imperative to take conscious steps to mitigate these biases.
+This can include diversifying the training data which is out of the scope of this paper.
+Figure 8: Upscaling: the results of our method (top row) may be further improved using an upscaler (bottom row) [1].
 5 Conclusion
-In this paper, we have demonstrated how full-body human generation and subsequent inpainting can be made context-aware, i.e., our generated humans are in sync with the environment when their clothing semantics and overall color harmonization are considered.
+We have demonstrated how full-body human generation and subsequent inpainting can be made context-aware, i.e., our generated humans are in sync with the environment when their clothing semantics and overall color harmonization are considered.
 Our main contribution is conditioning the image generation process on scene context to generate environment-specific people in clothing that is suitable for inpainting into a given environment.
-Our results demonstrate that our method produces synthetic imagery of plausible contextualized people in scenes, outperforming state-of-the-art methods on the task of contextual full-body generation.
-Acknowledgement: This project has received funding from the EU Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No.
-860768.
-MJB has received research gift funds from Adobe, Intel, Nvidia, Meta/Facebook, and Amazon.
-MJB has financial interests in Amazon, Datagen Technologies, and Meshcapade GmbH.
-While MJB is a consultant for Meshcapade, his research in this project was performed solely at, and funded solely by, the Max Planck Society.
-Appendix
-Appendix A Model architecture
-Scene representation.
-We extract the contextual scene representation vectors from the latent space of a variational auto-encoder (VAE).
-Here, the encoder consists of $5$ blocks of layers, where each block is comprised of one convolutional, batch normalization, and a leaky ReLU layer.
-The bottleneck layer is $512$-dimensional, and it contains our context code that is used as a conditioning in the StyleGAN generators for human parsing maps.
-The decoder mirrors the encoder, and its $5$ blocks are comprised of upconvolution, batch normalization, and a leaky ReLu layer.
-In the last block, we add an additional convolution and a tanh activation.
-The architecture of the remaining modules follows the guidelines defined in Style-GAN2 [22], Stable Diffusion LDM [41], and ControlNet [52]. Code will be released.
-Appendix B Training details
-Scene representation.
-We train the scene VAE on the AHP dataset [57] with a batch size of $64$ for approximately $500$ epochs and with a learning rate of $0.0002$.
-We use Adam optimizer with betas set to $0.5$ and $0.9$.
-The weights for the reconstruction and KL divergence loss terms are set to $\lambda_{1}=1$ and $\lambda_{2}=10$ (see Equation 1.
-in the main paper).
-After the training procedure has been completed, we freeze the VAE’s parameters and use its latent space as the context embedding to condition the contextual human parsing map generators.
-The training procedure for the remaining modules follows the guidelines defined in Style-GAN2 [22], Stable Diffusion LDM [41], and ControlNet [52]. Code will be released.
-Figure 9: Additional results for different contexts.
-Appendix C Additional Results
-Figure 9 contains more results for various environments.
+Our results demonstrate that our method produces plausible contextualized people in scenes, outperforming state-of-the-art methods on the task of contextual full-body generation.
+Acknowledgements: We thank Soubhik Sanyal for his assistance in writing, and Benjamin Pellkofer for IT support.
+This work has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska Curie grant agreement No 860768 (CLIPE project).